OpenStudy (anonymous):

How do i solve 2(2x)^(1/3) +1=5

OpenStudy (anonymous):

subtract 1 divide by 2 cube

OpenStudy (anonymous):

2x^1/3=2?

OpenStudy (anonymous):

\[2(2x)^{\frac{1}{3}}+1=5\] \[2(2x)^{\frac{1}{3}}=4\] \[2(2x)^{\frac{1}{3}}=2\] \[2x=8\]

OpenStudy (anonymous):

damn typo third line should be \[(2x)^{\frac{1}{3}}=2\]

OpenStudy (anonymous):

how do i handle the exponential 1/3?

OpenStudy (ajprincess):